### v1

@@ Abstract
Wildfires pose a major threat to Portugal, with an average of over 115,000 hectares burned annually in the 45-year period of 1980-2024. Beyond a high number of ignitions, the country has experienced devastating mega-fires, such as those in 2017. Accurate forecasting of wildfire occurrence and burned areas is therefore essential for effective firefighting resource allocation and emergency preparedness. In this study, we present a novel two-stage ensemble approach that extends the widely used latent Gaussian modelling framework with the integrated nested Laplace approximation (INLA) for spatio-temporal wildfire forecasting. The first stage uses XGBoost, a gradient boosting model, to identify wildfire patterns from environmental covariates and historical fire records, producing one-month-ahead point forecasts for fire counts and burned area. These predictions are then incorporated as external covariates in a latent Gaussian model, which includes additional spatiotemporal random effects to produce the final probabilistic forecasts of monthly total fire counts and burned area at the council level. To effectively model both moderate and extreme wildfire events, we implement the extended generalised Pareto (eGP) likelihood (a sub-asymptotic distribution) within the INLA framework. We also develop and discuss penalised complexity priors (PC-priors) for the eGP parameters and provide a comprehensive comparison of the eGP likelihood against other commonly employed distributions in environmental modelling, such as the Gamma and Weibull distributions. The proposed framework addresses the challenge of accessing future environmental covariates, which are typically unavailable at prediction time, and demonstrates strong performance in one-month-ahead wildfire forecasting.
@@ 1 Introduction
Portugal is among the countries most severely affected by wildfires, due to its mild and humid winters followed by hot and dry summers, strong winds, and extensive areas with forests and shrublands. According to official records, in 2017 alone, more than 21,000 fire ignitions were recorded, resulting in over 540,000 hectares burned (DaCamara,, 2024). In addition to the high number of fire events and extent of burnt area, Portugal regularly experiences ‘mega-fires’, i.e., individual wildfires with extreme consequences. For instance, the deadly wildfires of 17 June 2017 claimed at least 66 lives and affected more than 220,000 hectares in 24 hours. These types of events result not only in economic losses and human casualties but also cause substantial environmental damage, including widespread deforestation. Consequently, the development of an effective wildfire forecasting system is crucial for enabling early warnings and better allocation of firefighting resources (DaCamara et al.,, 2018).
@@
The increasing demand for accurate wildfire modelling has led to a wide range of studies employing statistical and machine learning methods. These approaches can be broadly categorised based on how they represent wildfires. One common approach models wildfire as an event occurring at an ignition point, optionally with an associated burnt area. This perspective motivates point process models for fire ignitions (Xu and Schoenberg,, 2011; Gabriel et al.,, 2017; Opitz et al.,, 2020; Woolford et al.,, 2021) and marked point processes where the burnt area serves as the mark (Tonini et al.,, 2017; Xi et al.,, 2019; Koh et al.,, 2023; de Rivera et al.,, 2024; Duvsten Östin, Hanna and Gasslander, Tilda,, 2025). Alternatively, ignition events and associated burnt areas can be aggregated over spatial partitions, with models targeting either the total burnt area or both fire count and burnt area per unit, leading to areal modelling approaches (Opitz,, 2023; Cisneros et al.,, 2024; Lawler and Shaby,, 2024).
@@
To capture the risk of extremely large burnt areas, extreme value theory (EVT) is often employed to estimate high quantiles of the burnt area distribution. The Generalised Pareto Distribution (GPD) is a widely used EVT-based model for peak-over-threshold methods and has been applied to model exceedance probabilities (Pimont et al.,, 2021; Richards et al.,, 2023; Koh et al.,, 2023). When modelling the entire distribution of burnt areas, however, GPD-based methods require an auxiliary distribution for values below the threshold, which can lead to discontinuities in the likelihood. Recent advances have introduced sub-asymptotic distributions, which offer continuous density, flexible tail behaviour, and theoretical justification within EVT (Papastathopoulos and Tawn,, 2013; Naveau et al.,, 2016). These distributions have seen successful applications in domains such as precipitation (Naveau et al.,, 2016), landslides (Yadav et al.,, 2021), and wildfires (Cisneros et al.,, 2024; Lawler and Shaby,, 2024).
@@
Bayesian hierarchical models with latent Gaussian fields are a popular framework for high-dimensional spatial and spatio-temporal modelling (Opitz,, 2017). While inference is traditionally performed via Markov Chain Monte Carlo (MCMC), the integrated nested Laplace approximation (INLA) offers a faster and accurate alternative for posterior approximation, especially in space-time applications (Rue et al.,, 2009). Several studies have applied INLA-based frameworks to wildfire modelling (Gabriel et al.,, 2017; Opitz et al.,, 2020; Koh et al.,, 2023). Machine learning methods have also been explored, including tree-based models (Koh,, 2023; Cisneros et al.,, 2023) and neural networks (Richards and Huser,, 2022; Richards et al.,, 2023; Cisneros et al.,, 2024).
@@
Despite the breadth of existing research, practical wildfire forecasting methods remain limited. The development and spread of wildfires are strongly influenced by environmental factors such as humidity, temperature, wind, and vegetation type. Incorporating such information is essential for improving predictive accuracy. However, most current spatio-temporal frameworks (e.g. Koh et al.,, 2023; Cisneros et al.,, 2024) are designed for retrospective analysis, using covariates observed at time $t$ to predict wildfires also occurring at time $t$. This setup limits their use in real-time forecasting, as it assumes access to future covariates that would not be available at the times for which predictions are to be made. Furthermore, within the popular latent Gaussian modelling framework, the additive structure and the practical constraints on the number of hyperparameters restrict the inclusion of multiple covariates in INLA-based models. As a result, studies often rely on a small number of representative variables, such as the Fire Weather Index (FWI), an index developed by the Canadian Forestry Service that has proven to be an especially suitable indicator of meteorological fire danger in Mediterranean ecosystems (DaCamara et al.,, 2014; Pinto et al.,, 2018; Nunes et al.,, 2023)
@@
In this work, we aim to address the twin challenges of acquiring future covariates and the limited capacity of the INLA framework to accommodate numerous predictors. We propose a two-stage, interpretable modelling framework that relies entirely on readily available reanalysis data for probabilistic forecasting, leveraging both machine learning and INLA. In the first stage, we train a tree-based ensemble model, specifically, XGBoost (Chen and Guestrin,, 2016), on a window-based dataset, incorporating environmental covariates and historical wildfire data up to time $t$, with the target being wildfire activity at time $t+1$. This model learns patterns from historical data to produce a point forecast for the next time step. In the second stage, the XGBoost forecast is used as a synthetic future covariate, combined with spatial and temporal Gaussian effects in a latent Gaussian model estimated via INLA, yielding posterior predictive distributions.
The XGBoost model effectively encodes the information from all available covariates into a single, most informative predictor for Portuguese wildfires. This alleviates the need for future environmental covariates and circumvents the limitations of INLA in handling many covariates. Meanwhile, INLA provides a principled framework for uncertainty quantification through the posterior predictive distribution. By integrating the strengths of both approaches, this stacked modelling strategy has the potential to enhance predictive performance (Wolpert,, 1992).
@@
Additionally, we contribute to the integration of the extended Generalised Pareto (eGP) likelihood (Naveau et al.,, 2016) within the INLA framework. The eGP distribution belongs to the family of sub-asymptotic models capable of jointly modelling the bulk and tail of the data. We use this distribution to model burnt area data, thereby capturing the moderate and extreme fires simultaneously in a continuous manner. As part of our implementation, we derive and incorporate penalised complexity (PC) priors with a closed-form expression for the two shape parameters of the eGP distribution.
@@ 2.1 Wildfire Data Scope
Administrative metadata, including council and district information, is mapped to each fire record. Fires are then aggregated to the council-month level to facilitate areal modelling. This choice is motivated by two factors: (1) The recorded ignition coordinates lack high spatial precision, and repeated wildfires are often observed at the same location, making a regional modelling unit such as the council more appropriate. (2) Council-level predictions are more interpretable and actionable for policymakers than point-level estimates. As such, modelling directly on aggregated data is preferred over integrating point process models post hoc. The temporal aggregation to a monthly resolution is a deliberate balance between computational feasibility and practical relevance. Modelling at higher temporal resolution (e.g. daily or hourly) over a 13-year span would be computationally intensive and likely require oversimplified spatio-temporal structures, potentially compromising model accuracy. Monthly aggregation allows for a more complex spatio-temporal modelling without excessive computational burden (Krainski et al.,, 2018).
@@
Figure 1: Histograms of fire count and burnt area at the council-month level, highlighting the prevalence of zeros.
Figure 2: Top: Average council-level fire count and burnt area. Bottom: Monthly average fire count and burnt area across all councils. Spatial and seasonal variation is evident.
@@ 3.1 Overview
We propose a two-stage modelling framework to provide probabilistic forecasts of wildfires at the monthly council level. In the first stage, we employ the XGBoost algorithm to integrate historical wildfire activity data and current environmental covariates to generate one-month-ahead forecasts of wildfire count and burnt area for each council. These forecasts are then passed to a latent Gaussian model estimated via INLA, which incorporates spatio-temporal dependencies through council adjacency and temporal encoding. The INLA framework outputs full posterior predictive distributions for fire presence, fire count, and burnt area. To model the heavy-tailed behaviour of burnt areas, we implement the extended generalised Pareto (eGP) likelihood in INLA, which blends a Gamma-like left tail with a Pareto-like right tail, and is characterised by two shape parameters and a scale parameter (see Section 3.3.2 for details). This two-stage approach can be viewed as a form of model stacking (Wolpert,, 1992), a type of ensemble learning where the predictions of one model are used as input features for another. Figure 3 provides a high-level overview of the proposed modelling pipeline.
@@
The combination of XGBoost and INLA utilises the strengths of each framework while mitigating their individual limitations. XGBoost is particularly effective in modelling complex, nonlinear interactions among environmental and historical variables. By incorporating past wildfire activity in an autoregressive fashion, XGBoost can compensate for the potential loss of information due to the spatial smoothing inherent in gridded environmental data. However, a key drawback of XGBoost is its lack of native uncertainty quantification. This is addressed in the second stage of our framework, where the INLA-based Bayesian hierarchical model provides full posterior distributions, allowing for straightforward uncertainty quantification.
@@ 3.2 Stage I: XGBoost
XGBoost is a scalable and efficient gradient boosting framework (Friedman,, 2001) designed for structured data. As an ensemble method, it constructs a strong predictive model by sequentially combining multiple tree models, each of which serves as a weak learner. The fundamental idea is to improve model performance by learning residual patterns from previous models in an additive and iterative fashion.
@@
$\mathcal{L}^{(m)}=\sum_{i}\ell(y_{i},\widehat{y}_{i}^{(m-1)}+f_{m}(\boldsymbol{x}_{i}))+\Omega(f_{m}),$
[…]
$\mathcal{L}^{(m)}\approx\sum_{i}\left[\ell(y_{i},\widehat{y}_{i}^{(m-1)})+g_{i}f_{m}(\boldsymbol{x}_{i})+\frac{1}{2}h_{i}f_{m}^{2}(\boldsymbol{x}_{i})\right]+\gamma L+\frac{1}{2}\lambda\sum_{j=1}^{L}w_{j}^{2},$
@@
The choice of loss function $\ell$ is central to the performance of XGBoost and should align with the distributional properties of the response variable. In our case, fire count is a non-negative integer and is naturally modelled using a Poisson loss. By contrast, the burnt area is of a mixed type: it is continuous and positive when fires occur, but has a point mass at zero when no fire is observed. For this, we adopt the Tweedie deviance loss (Jørgensen,, 1987), which corresponds to a compound Poisson-Gamma distribution. This distribution models a sum of Gamma-distributed fire sizes conditional on a Poisson-distributed number of occurrences:
@@ 3.2.1 Window-Based Modelling
Tree-based models, including XGBoost, do not inherently account for sequential dependencies in time-series data. A naive implementation would treat each time point independently, using covariates $\boldsymbol{x}_{t}$ to predict the target $y_{t}$, thereby neglecting temporal autocorrelation and requiring future covariates for forecasting. Given our 1-month modelling granularity, it is challenging to obtain the future environmental covariates in such a large horizon.
To address this limitation, we adopt a window-based approach, a common practice in time-series forecasting for non-sequential models (Elsayed et al.,, 2021). For a given forecasting horizon of one month, we reformulate the data into a lagged autoregressive structure: each target $y_{t}$ is modelled as a function of covariates and wildfire records from previous time steps, such as $(\boldsymbol{x}_{t-1},y_{t-1},y_{t-2},\ldots,y_{t-w})$, where $w$ is the time window size. This configuration enables the model to learn temporal dependencies and make forecasts based on available historical wildfire data and covariates. It also helps mitigate the issue of smoothed covariate values, which are uninformative for local fire prediction, by incorporating recent wildfire history. Figure 4 contrasts the window-based modelling with the naive approach.
@@
Let $\widetilde{\boldsymbol{x}}_{s,t}^{C}$ and $\widetilde{\boldsymbol{x}}_{s,t}^{B}$ denote the complete feature vectors used to forecast fire count ($C$) and burnt area ($B$), respectively, at council $s$ and time $t$. Both vectors share the same covariates listed in A.1 and A.2, except $\widetilde{\boldsymbol{x}}_{s,t}^{C}$ includes only fire count-based autoregressive covariates, while $\widetilde{\boldsymbol{x}}_{s,t}^{B}$ includes only those based on burnt area. The forecasts for fire count and burnt area at council $s$ and time $t+1$ are then given by:

### v2

@@ Abstract
Wildfires pose a major threat to Portugal, with over 115,000 hectares burned annually on average during 1980-2024, and the country has faced devastating mega-fires such as those in 2017. Accurate forecasts of wildfire occurrence and burned area are therefore essential for firefighting resource allocation and emergency preparedness. In this study, we propose a novel two-stage ensemble that extends the widely used latent Gaussian modelling framework with integrated nested Laplace approximation (INLA) for spatio-temporal wildfire forecasting. Stage 1 applies a gradient boosting model (XGBoost) to environmental covariates and historical fire records to produce one-month-ahead point forecasts of fire counts and burned area. Stage 2 uses these predictions as external covariates in a latent Gaussian model with additional spatiotemporal random effects to generate probabilistic forecasts of monthly total fire counts and burned area at the council level. To capture both moderate and extreme events, we implement the extended generalised Pareto (eGP) likelihood (a sub-asymptotic distribution) within INLA, develop Penalised Complexity (PC) priors for its parameters, and compare the eGP likelihood with common alternatives (e.g., Gamma and Weibull). Our framework tackles the unavailability of future environmental covariates at prediction time and performs strongly for one-month-ahead forecasts.
@@ 1 Introduction
Portugal is among the countries most severely affected by wildfires, due to its mild and humid winters followed by hot and dry summers, strong winds, and extensive areas with forests and shrublands. According to official records, in 2017 alone, more than 21,000 fire ignitions were recorded, resulting in over 540,000 hectares burned (DaCamara, 2024). In addition to the high number of fire events and extent of burnt area, Portugal regularly experiences mega-fires, i.e., individual wildfires with extreme consequences. For instance, the deadly wildfires of 17 June 2017 claimed at least 66 lives and affected more than 220,000 hectares in 24 hours. These types of events result not only in economic losses and human casualties but also cause substantial environmental damage, including widespread deforestation. Consequently, the development of an effective wildfire forecasting system is crucial for enabling early warnings and better allocation of firefighting resources (DaCamara et al., 2018).
@@
The increasing demand for accurate wildfire modelling has led to a wide range of studies employing statistical and machine learning methods. These approaches can be broadly categorised based on how they represent wildfires. One common approach models wildfire as an event occurring at an ignition point, optionally with an associated burnt area. This perspective motivates point process models for fire ignitions (Xu and Schoenberg, 2011; Gabriel et al., 2017; Opitz et al., 2020; Woolford et al., 2021) and marked point processes where the burnt area serves as the mark (Tonini et al., 2017; Xi et al., 2019; Koh et al., 2023; de Rivera et al., 2024; Duvsten Östin, Hanna and Gasslander, Tilda, 2025). Alternatively, ignition events and associated burnt areas can be aggregated over spatial partitions, with models targeting either the total burnt area or both fire count and burnt area per unit, leading to areal modelling approaches (Opitz, 2023; Cisneros et al., 2024; Lawler and Shaby, 2024).
@@
To capture the risk of extremely large burnt areas, extreme value theory (EVT) is often employed to estimate high quantiles of the burnt area distribution. The Generalised Pareto Distribution (GPD) is a widely used EVT-based model for peak-over-threshold methods and has been applied to model exceedance probabilities (Pimont et al., 2021; Richards et al., 2023; Koh et al., 2023). When modelling the entire distribution of burnt areas, however, GPD-based methods require an auxiliary distribution for values below the threshold, which can lead to discontinuities in the likelihood. Recent advances have introduced sub-asymptotic distributions, which offer continuous density, flexible tail behaviour, and theoretical justification within EVT (Papastathopoulos and Tawn, 2013; Naveau et al., 2016). These distributions have seen successful applications in domains such as precipitation (Naveau et al., 2016), landslides (Yadav et al., 2021), and wildfires (Cisneros et al., 2024; Lawler and Shaby, 2024).
@@
Bayesian hierarchical models with latent Gaussian fields are a popular framework for high-dimensional spatial and spatio-temporal modelling (Opitz, 2017). While inference is traditionally performed via Markov Chain Monte Carlo (MCMC), the integrated nested Laplace approximation (INLA) offers a faster and accurate alternative for posterior approximation, especially in space-time applications (Rue et al., 2009). Several studies have applied INLA-based frameworks to wildfire modelling (Gabriel et al., 2017; Opitz et al., 2020; Koh et al., 2023). Machine learning methods have also been explored, including tree-based models (Koh, 2023; Cisneros et al., 2023) and neural networks (Richards and Huser, 2022; Richards et al., 2023; Cisneros et al., 2024).
@@
Despite the breadth of existing research, practical wildfire forecasting methods remain limited. The development and spread of wildfires are strongly influenced by environmental factors such as humidity, temperature, wind, and vegetation type. Incorporating such information is essential for improving predictive accuracy. However, most current spatio-temporal frameworks (e.g. Koh et al., 2023; Cisneros et al., 2024) are designed for retrospective analysis, using covariates observed at time $t+1$ to predict wildfires also occurring at time $t+1$. This setup limits their use in real-time forecasting, as it assumes access to future covariates that would not be available at the times for which predictions are to be made. Furthermore, within the popular latent Gaussian modelling framework, the additive structure and the practical constraints on the number of hyperparameters restrict the inclusion of multiple covariates in INLA-based models. As a result, studies often rely on a small number of representative variables, such as the Fire Weather Index (FWI), an index developed by the Canadian Forestry Service that has proven to be an especially suitable indicator of meteorological fire danger in Mediterranean ecosystems (DaCamara et al., 2014; Pinto et al., 2018; Nunes et al., 2023)
@@
In this work, we aim to address the twin challenges of acquiring future covariates and the limited capacity of the INLA framework to accommodate numerous predictors. We propose a two-stage, interpretable modelling framework that relies entirely on readily available reanalysis data for probabilistic forecasting, leveraging both machine learning and INLA. In the first stage, we train a tree-based ensemble model, specifically, XGBoost (Chen and Guestrin, 2016), on a window-based dataset, incorporating environmental covariates and historical wildfire data up to time $t$, with the target being wildfire activity at time $t+1$. This model learns patterns from historical data to produce a point forecast for the next time step. In the second stage, the XGBoost forecast is used as a synthetic future covariate, combined with spatial and temporal Gaussian effects in a latent Gaussian model estimated via INLA, yielding posterior predictive distributions.
The XGBoost model effectively encodes the information from all available covariates into a single, most informative predictor for Portuguese wildfires. This strategy reduces the reliance on future environmental covariates, circumvents the limitations of INLA in handling a large number of predictors, and enhances the INLA model’s ability to capture complex interactions among covariates. Meanwhile, INLA provides a principled Bayesian framework for uncertainty quantification through the posterior predictive distribution. This hybrid modelling framework can be viewed as a stacking strategy (Wolpert, 1992), which has been shown to yield substantial improvements in predictive performance (Van der Laan et al., 2007). In contrast to other hybrid approaches for spatial and spatio-temporal modelling that require specialised mappings between spatial processes or covariance structures (Wikle and Zammit-Mangion, 2023), our approach minimises modifications to the underlying spatial structure, thereby preserving interpretability and facilitating its application to other spatial or spatio-temporal modelling tasks within the INLA framework. Our proposed method is also closely related to residual-learning approaches (MacBride et al., 2025), but does not rely on iterative updates between the first-stage and second-stage models. In our setting, such iterative training would be computationally expensive due to the inclusion of complex spatio-temporal random effects in the INLA model, but the potential gains from iterative residual learning are likely to be marginal.
@@
Additionally, we contribute to the integration of the extended Generalised Pareto (eGP) likelihood (Naveau et al., 2016) within the INLA framework. The eGP distribution belongs to the family of sub-asymptotic models capable of jointly modelling the bulk and tail of the data. We use this distribution to model burnt area data, thereby capturing the moderate and extreme fires simultaneously in a continuous manner. As part of our implementation, we derive and incorporate penalised complexity (PC) priors with a closed-form expression for the two shape parameters of the eGP distribution.
@@ 2.1 Wildfire Data Scope
Administrative metadata, including council and district information, is mapped to each fire record. Fires are then aggregated to the council-month level to facilitate areal modelling. This choice is motivated by two factors: (1) The recorded ignition coordinates lack high spatial precision, and repeated wildfires are often observed at the same location, making a regional modelling unit such as the council more appropriate. (2) Council-level predictions are more interpretable and actionable for policymakers than point-level estimates. As such, modelling directly on aggregated data is preferred over integrating point process models post hoc. The temporal aggregation to a monthly resolution is a deliberate balance between computational feasibility and practical relevance. Modelling at higher temporal resolution (e.g. daily or hourly) over a 13-year span would be computationally intensive and likely require oversimplified spatio-temporal structures, potentially compromising model accuracy. Monthly aggregation allows for a more complex spatio-temporal modelling without excessive computational burden (Krainski et al., 2018).
@@
Figure 1: Histograms of fire count and burnt area at the council-month level, highlighting the prevalence of zeros. Quantities are rescaled for visualisation purposes using a square-root transformation for fire counts and a logarithmic transformation for burned area.
Figure 2: Top: Average council-level fire count and burnt area. Bottom: Monthly average fire count and burnt area across all councils. Spatial and seasonal variation is evident. Quantities are rescaled for visualisation purposes using a square-root transformation for fire counts and a logarithmic transformation for burned area.
@@ 3.1 Overview
We propose a two-stage modelling framework to provide probabilistic forecasts of wildfires at the monthly council level. In the first stage, we employ the XGBoost algorithm to integrate historical wildfire activity data and current environmental covariates to generate one-month-ahead forecasts of wildfire count and burnt area for each council. These forecasts are then passed to a latent Gaussian model estimated via INLA, which incorporates spatio-temporal dependencies through council adjacency and temporal encoding. The INLA framework outputs full posterior predictive distributions for fire presence, fire count, and burnt area. To model the heavy-tailed behaviour of burnt areas, we implement the extended generalised Pareto (eGP) likelihood in INLA, which blends a Gamma-like left tail with a Pareto-like right tail, and is characterised by two shape parameters and a scale parameter (see Section 3.3.2 for details). This two-stage approach can be viewed as a form of model stacking (Wolpert, 1992), a type of ensemble learning where the predictions of one model are used as input features for another. Figure 3 provides a high-level overview of the proposed modelling pipeline.
@@
The combination of XGBoost and INLA utilises the strengths of each framework while mitigating their individual limitations. XGBoost is particularly effective in modelling complex, nonlinear interactions among environmental and historical variables. By incorporating past wildfire activity in an autoregressive manner, the two XGBoost models, one for fire count and one for burnt area, can partially recover information that may be lost due to the spatial smoothing inherent in gridded environmental data. However, a key drawback of XGBoost is its lack of native uncertainty quantification. This is addressed in the second stage of our framework, where the INLA-based Bayesian hierarchical model provides full posterior distributions, allowing for straightforward uncertainty quantification.
@@ 3.2 Stage I: XGBoost
XGBoost is a scalable and efficient gradient boosting framework (Friedman, 2001) designed for structured data. As an ensemble method, it constructs a strong predictive model by sequentially combining multiple tree models, each of which serves as a weak learner. The fundamental idea is to improve model performance by learning residual patterns from previous models in an additive and iterative fashion.
@@
$\mathcal{L}^{(m)}=\sum_{i=1}^{n}\ell(y_{i},\widehat{y}_{i}^{(m-1)}+f_{m}(\boldsymbol{x}_{i}))+\Omega(f_{m}),$
where $\ell(\cdot)$ is a differentiable convex loss function, $\widehat{y}_{i}^{(m-1)}=\sum_{j=1}^{m-1}f_{j}(\boldsymbol{x}_{i})$ is the prediction from the ensemble up to iteration m1m-1, and $\Omega(f_{m})=\gamma L+\frac{1}{2}\lambda\|\boldsymbol{w}\|^{2}$ penalises the model complexity through the number of leaves $L$ and the leaf weights $\boldsymbol{w}$. To improve the optimisation efficiency, the loss is approximated using a second-order Taylor expansion (Chen and Guestrin, 2016):
$\mathcal{L}^{(m)}\approx\sum_{i=1}^{n}\left[\ell(y_{i},\widehat{y}_{i}^{(m-1)})+g_{i}f_{m}(\boldsymbol{x}_{i})+\frac{1}{2}h_{i}f_{m}^{2}(\boldsymbol{x}_{i})\right]+\gamma L+\frac{1}{2}\lambda\sum_{j=1}^{L}w_{j}^{2},$
@@
The choice of loss function $\ell$ is central to the performance of XGBoost and should align with the distributional properties of the response variable. In our case, fire count is a non-negative integer and is naturally modelled using a Poisson loss. By contrast, the burnt area is of a mixed type: it is continuous and positive when fires occur, but has a point mass at zero when no fire is observed. For this, we adopt the Tweedie deviance loss (Jørgensen, 1987), which corresponds to a compound Poisson-Gamma distribution. This distribution models a sum of Gamma-distributed fire sizes conditional on a Poisson-distributed number of occurrences:
@@ 3.2.1 Window-Based Modelling
Tree-based models, including XGBoost, do not inherently account for sequential dependencies in time-series data. A naive implementation would treat each time point independently, using covariates $\boldsymbol{x}_{t+1}$ to predict the target $y_{t+1}$, thereby neglecting temporal autocorrelation and requiring future covariates for forecasting. Given our 1-month modelling granularity, it is challenging to obtain the future environmental covariates over such a large horizon.
To address this limitation, we adopt a window-based approach, a common practice in time-series forecasting for non-sequential models (Elsayed et al., 2021). For a given forecasting horizon of one month, we reformulate the data into a lagged autoregressive structure: each target $y_{t+1}$ is modelled as a function of covariates and wildfire records from previous time steps, such as $(\boldsymbol{x}_{t},y_{t},y_{t-1},\ldots,y_{t-w+1})$, where $w$ is the time window size. This configuration enables the model to learn temporal dependencies and make forecasts based on available historical wildfire data and covariates. It also helps mitigate the issue of smoothed covariate values, which are uninformative for local fire prediction, by incorporating recent wildfire history. Figure 4 contrasts the window-based modelling with the naive approach.
@@
Let $\widetilde{\boldsymbol{x}}_{s,t}^{C}$ and $\widetilde{\boldsymbol{x}}_{s,t}^{B}$ denote the complete feature vectors used to forecast fire count ($C$) and burnt area ($B$), respectively, at council $s$ and time $t$. Both vectors share the same covariates listed in Table A.1 and A.2, except $\widetilde{\boldsymbol{x}}_{s,t}^{C}$ includes only fire count-based autoregressive covariates, while $\widetilde{\boldsymbol{x}}_{s,t}^{B}$ includes only those based on burnt area. The forecasts for fire count and burnt area at council $s$ and time $t+1$ are then given by: